Consider the composition of a 3-child family in which the children were born at different times but a girl is as likely as a boy at each birth. What is the probability that the oldest child and the youngest child are of the same gender?


1
Expert's answer
2022-03-21T06:49:08-0400

Let's denote B for the birth of the boy and G for the birth of the girl.


There 8 possible outcomes:


BBB, BBG, BGB, GBB, BGG, GBG, GGB, GGG


the oldest child and the youngest child are of the same gender: BGB, GBG


Therefore the probability that the oldest child and the youngest child are of the same gender is: 2/8=1/4=0.25